# Kirby diagrams, trisections and gems of PL 4-manifolds: relationships, results and open problems

### v2

@@ 3.2.3 Second indirect approach
([12])  Let $\Gamma\in G_{s}^{(4)}$ be a gem of a compact $4$-manifold $M$ with $\partial M\cong\#_{m}(\mathbb{S}^{1}\otimes\mathbb{S}^{2})$, $m\geq 0$. For each cyclic permutation $\varepsilon$ of $\Delta_{4}$, the triple $\mathcal{T}(\Gamma,\varepsilon)$ gives rise to a trisection of $\bar{M}\cong M\cup{\mathbb{Y}}_{m}^{(\sim)}$ by applying $k$ stabilizations, with $0\leq k\leq\rho_{\varepsilon}(\Gamma)-\rho_{\varepsilon}(\Gamma_{\hat{4}})$. Hence:
@@
Further, well-known combinatorial formulas in gem theory (see for example [9, Proposition 3.1]) enable to check that $g_{\varepsilon_{0}\varepsilon_{3}}-g_{\varepsilon_{0}\varepsilon_{3}4}=\rho_{\varepsilon}(\Gamma)-\rho_{\varepsilon}(\Gamma_{\hat{4}})$.
Hence, a gem-induced trisection of $M$ is obtained, with central surface of genus: $\rho_{\varepsilon}(\Gamma_{\hat{4}})+(\rho_{\varepsilon}(\Gamma)-\rho_{\varepsilon}(\Gamma_{\hat{4}}))$. Finally, the thesis concerning $\bar{M}$ follows, by possible application of Theorem 13 (in case $m>0$).
@@ 3.3.1 From Kirby diagrams to trisection diagrams via gems
In [28], Martini and Toriumi analyze the decomposition generated by their procedure to yield trisections from any gem, and point out that it gives rise to a collection of trisection diagrams of the appropriate genus, since they obtain a redundant amount of attaching curves of the 3-dimensional intersections: see [28, Section 4.6] for details.
By adapting the idea of Martini and Toriumi to the case of a gem $\Gamma$ satisfying condition (*) of Proposition 11 (and therefore admitting a gem-induced trisection $\mathcal{T}(\Gamma,\varepsilon)$), a trisection diagram can actually be identified on the central surface of the induced trisection, which turns out to coincide with the embedding surface $F_{\varepsilon}(\Gamma_{\hat{4}})$ of the graph without color 4. $F_{\varepsilon}(\Gamma_{\hat{4}})$ has genus $\rho_{\varepsilon}(\Gamma_{\hat{4}})$ in the orientable case and $2\rho_{\varepsilon}(\Gamma_{\hat{4}})$ in the non-orientable one, and the systems of curves drawn on it and constituting a trisection diagram are nothing but suitable bicolored cycles of the given gem.
@@
Then, for each cyclic permutation $\varepsilon=(\varepsilon_{0},\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},4)$ of $\Delta_{4}$, a trisection diagram for the associated closed $4$-manifold $\bar{M}$ is given by the following three systems of curves $\{\alpha,\beta,\gamma\}$ on $F_{\varepsilon}(\Gamma_{\hat{4}}):$
@@
Proof. According to Theorem 10, the gem-induced trisection $\mathcal{T}(\Gamma,\varepsilon)$ of $M$ is formed by the triple $(H_{0},H_{1},H_{2})$, where $H_{1}$ (resp. $H_{2}$) is a regular neighbourhood of $K_{\varepsilon_{0}\varepsilon_{2}}$ (resp. $K_{\varepsilon_{1}\varepsilon_{3}}$), while $H_{0}$ is either the $4$-disk obtained by coning over the triangulation $K(\Gamma_{\hat{4}})$ of the $3$-sphere (if $m=0$) or a collar of the triangulation $K(\Gamma_{\hat{4}})$ of $\partial M\cong\#_{m}(\mathbb{S}^{1}\otimes\mathbb{S}^{2})$ (if $m>0$).
@@
It is not difficult to check that, both in the case $m=0$ (where we define $\bar{\mathcal{T}}=\mathcal{T}(\Gamma,\varepsilon)$ for sake of notational simplicity) and in the case $m>0$, the trisection $\bar{\mathcal{T}}$ is completely identified by the attachments on $\Sigma$ of $H_{01}=H_{0}\cap H_{1}=\bar{H}_{0}\cap H_{1}$, $H_{02}=H_{0}\cap H_{2}=\bar{H}_{0}\cap H_{2}$ and $H_{12}=H_{1}\cap H_{2}$ (which are 3-dimensional handlebodies by hypothesis). Now, as pointed out by Martini and Toriumi [^12] See [28, Paragraph 4.6], and in particular the caption of Figure 17., the quadrangulation of the central surface $\Sigma$ given by the union of the “orange” squares in Figure 15 is dual to the cellular subdivision of the embedding surface $F_{\varepsilon}(\Gamma_{\hat{4}})$, given by $\Gamma_{\hat{4}}$ itself. Hence, $\Sigma$ and $F_{\varepsilon}(\Gamma_{\hat{4}})$ can be identified.
Moreover, each $\{\varepsilon_{0},\varepsilon_{2}\}$-colored (resp. $\{\varepsilon_{1},\varepsilon_{3}\}$-colored) cycle of $\Gamma$, that is embedded in $F_{\varepsilon}(\Gamma_{\hat{4}})$, is dual in $K(\Gamma_{\hat{4}})$ to a red (resp. green) edge to which a prism constituting $H_{02}$ (resp. $H_{01}$) collapses (see Figure 15 and [28, Paragraph 4.6]); further, the union of such edges, that is a spine of $H_{02}$ (resp. $H_{01}$), is isomorphic to $K_{\varepsilon_{1}\varepsilon_{3}}$ (resp. $K_{\varepsilon_{0}\varepsilon_{2}}$). Hence, the statement regarding $\alpha$ and $\beta$ curves easily follows, by noting that a system of independent meridian curves for $H_{02}$ (resp. $H_{01}$) can be obtained from the set of $\{\varepsilon_{0},\varepsilon_{2}\}$-colored cycles (resp. $\{\varepsilon_{1},\varepsilon_{3}\}$-colored cycles) by removing those corresponding to a maximal tree of $K_{\varepsilon_{1}\varepsilon_{3}}$ (resp. $K_{\varepsilon_{0}\varepsilon_{2}}$).
@@
Moreover, by hypothesis, $Q(\Gamma,\varepsilon)$ further collapses to a 1-dimensional complex via the collapsing sequence given by the ordering $(e_{1},\dots,e_{p})$ of $4$-colored edges satisfying condition (*), as described in the proof of Proposition 11. Hence, a spine of $H_{12}$ consists of the set of edges of $Q(\Gamma,\varepsilon)$ remaining at the end of the collapsing sequence, and a system of (independent) curves realizing the attachment of $H_{12}$ can be obtained from this set of edges by shrinking to a point the ones belonging to a maximal tree. The statement regarding the $\gamma$ curves follows, by projecting each element of the above system of curves (corresponding to a $\{c,4\}$-cycle of $\Gamma$) on $F_{\varepsilon}(\Gamma_{\hat{4}})$: in fact, while all $c$-colored edges ($c\in\Delta_{3}$) obviously embed into $F_{\varepsilon}(\Gamma_{\hat{4}})$, the ordering (*) enables to project on $F_{\varepsilon}(\Gamma_{\hat{4}})$ also the 4-colored edge $e_{j}$, for increasing values of $j\in\{1,\ldots,p\}$, via the bicolored cycle containing it whose edges different from $e_{j}$ already embed in $F_{\varepsilon}(\Gamma_{\hat{4}})$.
@@
    the $\beta$ curves are all $\{0,3\}$-colored cycles of $\Gamma(L,d)$ (resp. $\tilde{\Omega}(L,d)$), but $12l-$ 1 corresponding to a maximal tree of the 1-dimensional subcomplex $K_{12}$ generated by the vertices labelled $1$ and $2$;
@@
    first, in any order, all $4$-colored edges belonging to $\{1,4\}$-colored cycles of length two, together with the $l$ $4$-colored edges between the vertices $P_{0}^{(i)}$ and $P_{1}^{(i)}$, $\forall i\in\{1,2,\dots,l\}$ (which belong to $\{0,4\}$-colored cycles of length two);
    then, the $l$ $4$-colored edges between the vertices $P_{4}^{(i)}$ and $P_{5}^{(i)}$, $\forall i\in\{1,2,\dots,l\};$
    finally, the $l$ $4$-colored edges between the vertices $P_{2}^{(i)}$ and $P_{3}^{(i)}$, $\forall i\in\{1,2,\dots,l\}.$
@@
    $(\Sigma;\alpha,\gamma)$ and $(\Sigma;\beta,\gamma)$ are (genus $g$) Heegaard diagrams of #$\#_{k_{1}}(\mathbb{S}^{1}\times\mathbb{S}^{2})$ and #$\#_{k_{2}}(\mathbb{S}^{1}\times\mathbb{S}^{2})$ respectively, with $0\leq k_{i}\leq g$, for $i=1,2$;
@@
Then, for any cyclic permutation $\varepsilon$ of $\Delta_{4}$, a G-trisection diagram for $M$ is given by the following three system of curves $\{\alpha,\beta,\gamma\}$ on $F_{\varepsilon}(\Gamma_{\hat{4}})$:
@@
    the $\beta$ curves are all $\{0,3\}$-colored cycles of $\Gamma(L,d)$ (resp. $\tilde{\Omega}(L,d)$), but $12l-$ 1 corresponding to a maximal tree of the 1-dimensional subcomplex $K_{12}$ generated by the vertices labelled $1$ and $2$;
@@ 4 Open problems and trends
Another possible improvement would be to study how to obtain a gem of a closed 4-manifold, starting from a gem of the compact one consisting only of handles up to index two, even if no $\rho_{3}$- or $\rho_{4}$-pair appears in the latter (as assumed in Proposition 9). The problem is obviously a difficult one, since it involves the combinatorial recognition of #$\#_{m}(\mathbb{S}^{1}\times\mathbb{S}^{2})$ ($m>0$) in 3-dimensional topology, but in specific situations it could possibly be faced, for example by detecting sequences of combinatorial moves yielding the desired $\rho_{3}$-pairs.
For example, the gem depicted in Figure 22 (representing the orientable genus one 4-dimensional handlebody) contains no $\rho_{3}$-pairs in its $\hat{4}$-residue; however, by a sequence of combinatorial moves not affecting the represented 4-manifold (whose result is the “cyclic” exchanging of the end-points of three 1-colored edges), a $\rho_{3}$-pair of color $1$ appears in the $\hat{4}$-residue, which is also a $\rho_{3}$-pair in the 5-colored graph: see Figure 23, where the edges cyclically exchanged are highlighted in yellow, while the $\rho_{3}$-pair is highlighted in blue. The $\rho_{3}$-pair switching yields the 5-colored graph of Figure 24, which represents the 4-sphere; since a sphere-bundle summand is “lost” in the switching (according to Proposition 9(ii)), it is confirmed that the associated closed 4-manifold $\bar{M}$ is $\mathbb{S}^{1}\times\mathbb{S}^{3}$.
Figure 22: A Kirby diagram $(L,d)$ (with two dotted components) and the associated 5-colored graph representing $\mathbb{Y}^{4}_{1}$
Figure 23: Another 5-colored graph (with a $\rho_{3}$-pair) representing $\mathbb{Y}^{4}_{1}$
Figure 24: A 5-colored graph representing $N^{4}$ ($N^{4}\cong\mathbb{S}^{4}$), so that $\bar{M}\ =\ (\mathbb{S}^{1}\times\mathbb{S}^{3})\#N^{4}$

### v3

@@ 3.2.3 Second indirect approach
[…]
@@
[…]
Hence, a gem-induced trisection of $M$ is obtained, with central surface of genus: $)\rho_{\varepsilon}(\Gamma_{\hat{4}})+(\rho_{\varepsilon}(\Gamma)-\rho_{\varepsilon}(\Gamma_{\hat{4}}$)).
Finally, the thesis concerning $\bar{M}$ follows, by possible application of Theorem 13 (in case $m>0$).
@@ 3.3.1 From Kirby diagrams to trisection diagrams via gems
In [29], Martini and Toriumi analyze the decomposition generated by their procedure to yield trisections from any gem, and point out that it gives rise to a collection of trisection diagrams of the appropriate genus, since they obtain a redundant amount of attaching curves of the 3-dimensional intersections: see [29, Section 4.6] for details.
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
    the $\beta$ curves are all $\{0,3\}$-colored cycles of $\Gamma(L,d)$ (resp. $\tilde{\Omega}(L,d)$), but 2l12l-1 corresponding to a maximal tree of the 1-dimensional subcomplex $K_{12}$ generated by the vertices labelled $1$ and $2$;
@@
    first, in any order, all $4$-colored edges belonging to $\{1,4\}$-colored cycles of length two, together with the $l$ 44-colored edges between the vertices $P_{0}^{(i)}$ and $P_{1}^{(i)}$, $\forall i\in\{1,2,\dots,l\}$ (which belong to $\{0,4\}$-colored cycles of length two);
    then, the $l$ 44-colored edges between the vertices $P_{4}^{(i)}$ and $P_{5}^{(i)}$, $\forall i\in\{1,2,\dots,l\};$
    finally, the $l$ 44-colored edges between the vertices $P_{2}^{(i)}$ and $P_{3}^{(i)}$, $\forall i\in\{1,2,\dots,l\}.$
@@
    $(\Sigma;\alpha,\gamma)$ and $(\Sigma;\beta,\gamma)$ are (genus $g$) Heegaard diagrams of $_{k_{1}}(\mathbb{S}^{1}\times\mathbb{S}^{2})$ and $_{k_{2}}(\mathbb{S}^{1}\times\mathbb{S}^{2})$ respectively, with $0\leq k_{i}\leq g$, for $i=1,2$;
@@
[…]
@@
    the $\beta$ curves are all $\{0,3\}$-colored cycles of $\Gamma(L,d)$ (resp. $\tilde{\Omega}(L,d)$), but 2l12l-1 corresponding to a maximal tree of the 1-dimensional subcomplex $K_{12}$ generated by the vertices labelled $1$ and $2$;
@@
Remark 4
In [12], by using stabilizations along 4-colored edges and the ordering involved in the proof of Theorem 15, both Propositions 21 and 23 are extended to gems that do not directly induce trisections. As a consequence, trisection diagrams can be obtained algorithmically from any gem of a closed 4-manifold. Detailed statements and proofs can be found in [12], while Figures 22, 23 and 24 show an example of the procedure, applied to the standard order ten crystallization $\Gamma$ of $\mathbb{S}^{1}\times\mathbb{S}^{3}$ depicted in Figure 22. If a stabilization along the 4-colored edge $\bar{e}$ is performed, all squares of $Q(\Gamma,\varepsilon)$ ($\varepsilon=(0,1,2,3,4)$) subsequently collapse and the obtained 1-dimensional complex, after shrinking to a point a maximal tree, corresponds to the $\{2,4\}$-colored cycle highlighted in red in Figure 23. Hence, this cycle identifies the (unique) curve of the system $\gamma$ in the associated trisection diagram of genus one, whose surface is obtained, via the stabilization, from the 2-sphere where $\Gamma_{\hat{4}}$ regularly embeds. On the other hand, the systems $\alpha$ and $\beta$ consist of the blue and green trivial curves respectively, which are meridian curves of the 1-handles involved in the stabilization. It is easy to check that the obtained trisection diagram exactly coincides with the standard one of $\mathbb{S}^{1}\times\mathbb{S}^{3}$ (see Figure 24).
Figure 22: the standard order ten gem of $\mathbb{S}^{1}\times\mathbb{S}^{3}$ with a stabilizing edge highlighted
Figure 23: the curves of the trisection diagram of $\mathbb{S}^{1}\times\mathbb{S}^{3}$ induced by the gem
Figure 24: the associated trisection diagram of $\mathbb{S}^{1}\times\mathbb{S}^{3}$
@@ 4 Open problems and trends
Another possible improvement would be to study how to obtain a gem of a closed 4-manifold, starting from a gem of the compact one consisting only of handles up to index two, even if no $\rho_{3}$- or $\rho_{4}$-pair appears in the latter (as assumed in Proposition 9). The problem is obviously a difficult one, since it involves the combinatorial recognition of $_{m}(\mathbb{S}^{1}\times\mathbb{S}^{2})$ ($m>0$) in 3-dimensional topology, but in specific situations it could possibly be faced, for example by detecting sequences of combinatorial moves yielding the desired $\rho_{3}$-pairs.
A particular case of this problem arises, for instance, in the construction of a gem for the double of a compact $4$-manifold. It is known that a Kirby diagram of the double can be obtained, from a Kirby diagram of the manifold itself, by adding a 0-framed meridian to each framed component (see [24]). Therefore, it may be worth investigating whether the particular configuration of the arising 5-colored graph allows to be modified in a standard way so as to give rise to $\rho_{3}$-pairs satisfying Proposition 9.